I am having trouble trying to set up this word problem Two times the number of dimes in the bank equals the number of dimes plus 30. How many dimes are there in the bank? Please help me set up the problem so I can solve it. Thank you i am so confused!

Answer by macston(5194)

.
D=number of dimes in bank
.
Two times the number of dimes in the bank=2D
Number of dimes in bank plus 30=D+30
.
Two times the number of dimes=number of dimes + 30:
2D=D+30 . Subtract D from each side.
D=30
.
ANSWER: There are 30 dimes in the bank.
.
CHECK:
2D=D+30
2(30)=30+30
60=60